OpenStudy (anonymous):

Find the area of a triangle with a base length of 3 units and a height of 4 units. 3 units 6 units 7 units 8 units

OpenStudy (anonymous):

PLEASE HALP!!!!

OpenStudy (anonymous):

@sweetburger @dan815 @ganeshie8

OpenStudy (anonymous):

area of a triangle is base times height divided by 2

OpenStudy (anonymous):

Ty

OpenStudy (anonymous):

3 times 4 =12/2 =6

OpenStudy (anonymous):

So 6

OpenStudy (owlcoffee):

The area of a triangle is found by the semiproduct between the base and the height: \[A_{\triangle}=\frac{ b.h }{ 2 }\] or... applied: \[A _{\triangle}=\frac{ 3.4 }{ 2 }\]

OpenStudy (owlcoffee):

Yes, "6" is correct.